John can paint a room in 10 hours by himself. If kelly helps, they can paint the same room in 6 hours. If kelly works alone, how long would it take her to paint the room?
Answer by TimothyLamb(4379)

---
rates:
j = 1/10
j + k = 1/6
---
j + k = 1/6
1/10 + k = 1/6
k = 1/6 - 1/10
k = 10/60 - 6/60
k = 4/60
---
time = 1/rate:
kelly's time = 60/4
kelly's time = 15 hours
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php